A machine costing $185,000 with a five-year life and $20,000 residual value was purchased January 2. Compute depreciation for each of the five years, using the double-declining-balance method.

1.   Year 1$
2.   Year 2$
3.   Year 3$
4.   Year 4$
5.   Year 5$
Answer #1

Rate of Depreciation-1/5 x 2-40% Depreciation BookValue Year 1 $185,000 x 0.40 - $74,000 $185,000 $74,000 $111,000 Year 2 S11

Please do let me know if you have any questions
Answer #2
[REMOVED]
Answer #3

SOLUTION :


Machine cost = 185000 ($)


Salvage value = 20000 ($) 


Useful life = 5 years.


Depreciation rate = 1/5 = 0.20 = 20% 


Double depreciation rate = 2 * 20 = 40% 



Year     Beginning value of BV      Depreciation amount    Closing BV

                         ($)                                      ($)                              ($)



1                185000                         185000*0.4 = 74000               111000


2                 111000                         111000*0.4 = 44400                66600


3                   66600                         66600*0.4  = 26640                39960


                39960                          39960*0.4 = 15984                 23976


5                  23976                                                3976                  20000



answered by: Tulsiram Garg